OpenStudy (anonymous):

Solve for x. 9(x – 2) = 5(2 + x)

OpenStudy (black_ninja123):

First you have to simplify both of the equations

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

9(x-2)=5(2+x) left side: distribute the 9 to the terms inside the parenthesis and the right side: distribute the 5 to the terms inside the parenthesis

OpenStudy (anonymous):

(9*x) - (9*2) = (5*2) + (5*x) 9x - 18 = 10 + 5x subtract 5x from both sides

OpenStudy (anonymous):

4x - 18 = 10 add 18 to both sides

OpenStudy (anonymous):

4x = 28 --- --- 4 4 x = ?

OpenStudy (anonymous):

divide both sides by 4

OpenStudy (anonymous):

u got it ?

OpenStudy (anonymous):

x=8 thank yall :)

OpenStudy (anonymous):

np =) any time

OpenStudy (black_ninja123):

X=7 not 8